OpenStudy (anonymous):

I need an answer as fast as possible with explanation. Jogging burns about 100 calories per mile. Elizabeth jogs at a rate of 4.5 miles per hour. How long will it take her to burn off the 1200-calories lunch she ate at burger heaver?

OpenStudy (anonymous):

During on hour she burns 4.5*100=450 Calories so time=1200/450=8/3 hours or160minutes

OpenStudy (anonymous):

ThankYou!

OpenStudy (anonymous):

My pleasure!